OpenStudy (anonymous):

What are the possible number of negative zeros of f(x) = 2x7 – 2x6 + 7x5 – 7x4 – 4x3 + 4x2 ?

jhonyy9 (jhonyy9):

sorry but i think that the zero is zero and is not negative zero or exist positive zero too ?

OpenStudy (anonymous):

I will take a guess: 2, x=-2i and x=-2^(-1/2)\[2 x^7-2 x^6+7 x^5-7 x^4-4 x^3+4 x^2=(x-1) x^2 \left(x^2+4\right) \left(2 x^2-1\right) \]